OpenStudy (anonymous):

Calculate the molar mass of CH4 if 279 ml of the gas measured at 31 deg C and 492 mmHg has a mass of 0.116 g.

OpenStudy (anonymous):

Can someone just confirm my answer? I'm getting 16.0 g/mol Medal and a big fat appreciation given of course

OpenStudy (anonymous):

m.mass= RT X mass / PV R=0.821 T=31+273=304 K P=492/760=0.647 atm V=0.279 L m=0.116 g Answer will be 16.3 g/mol

OpenStudy (anonymous):

ur ans is right

OpenStudy (anonymous):

Thank you!